OpenStudy (anonymous):

Choose the correct description of the graph of the inequality x - 4 5x. Open circle on -1, shading to the left. Closed circle on -1, shading to the left. Open circle on -1, shading to the right. Closed circle on -1, shading to the right.

OpenStudy (anonymous):

I really need help! Can someone please help me by explaining im not asking for the answer im asking for someone to explain it so that I can figure the rest of them for myself.

OpenStudy (anonymous):

can't because something is missing in your question

OpenStudy (anonymous):

"Choose the correct description of the graph of the inequality x - 4 5x."

OpenStudy (anonymous):

x-4>= 5x

OpenStudy (anonymous):

\[x-4\geq 5x\] subtract \(x\) an get \[-4\geq 4x\] divide by \(4\) to get \[-1\geq x\] or \(x\leq -1\)

OpenStudy (anonymous):

in english, \(x\) is less than or equal to minus one it will have a closed circle because of the "equal to" part

OpenStudy (anonymous):

and since numbers that are less than minus one are to the left of \(-1\) pick " Closed circle on -1, shading to the left"

OpenStudy (anonymous):

Thank you soo much!!! That really helped a lot because I have 20 questions asking me the same thing just different equations.